At the beginning of the year, Addison Company's assets are $188,000 and its equity is $141,000. During the year, assets increase $80,000 and liabilities increase $47,000. What is the equity at year-end?

Assets = Liabilities + Equity
Beginning $188,000 = + $141,000
Change 80,000 = 47,000 +
Ending = +
Answer #1

Answer:

equity at year-end= $174,000

Assets = Liabilities + Equity
A B C=A-B
1 Beginning          1,88,000 =            47,000 +       1,41,000
2 Change             80,000 =            47,000 +          33,000
3=1+2 Ending          2,68,000 =            94,000 +       1,74,000
1)47000=188000-141000
2) 33000=80000-47000